# FS-Net: Full Scale Network and Adaptive Threshold for Improving Extraction of Micro-Retinal Vessel Structures

### v3

@@ Abstract
Retinal vascular segmentation, a widely researched topic in biomedical image processing, aims to reduce the workload of ophthalmologists in treating and detecting retinal disorders. Segmenting retinal vessels presents unique challenges; previous techniques often failed to effectively segment branches and microvascular structures. Recent neural network approaches struggle to balance local and global properties and frequently miss tiny end vessels, hindering the achievement of desired results. To address these issues in retinal vessel segmentation, we propose a comprehensive micro-vessel extraction mechanism based on an encoder-decoder neural network architecture. This network includes residual, encoder booster, bottleneck enhancement, squeeze, and excitation building blocks. These components synergistically enhance feature extraction and improve the prediction accuracy of the segmentation map. Our solution has been evaluated using the DRIVE, CHASE-DB1, and STARE datasets, yielding competitive results compared to previous studies. Given its accurate and robust performance, the proposed approach is a solid candidate for being implemented in real-life diagnostic centers and aiding ophthalmologists.
@@
Highlights should contain 3 to 5 bullet points with a maximum of 125 characters per bullet point, including spaces. Only the core results of the paper should be covered. Please ensure each bullet point is less than 125 characters 1. Encoder booster and bottleneck enhancement blocks are introduced to reduce feature loss in the encoder-decoder architecture 2. An adaptive threshold algorithm is proposed to find the optimal threshold for pixelwise classification 3. Micro-retinal vessel extraction was improved significantly 4. A cross-training evaluation was performed between the DRIVE and STARE datasets
@@ 1 Introduction
Medical images are crucial for diagnoses cavaro2010diagnostic , providing valuable information for medical experts to perform necessary procedures. However, these images can be noisy and difficult to interpret, particularly at the microlevel. This challenge is particularly noticeable in retinal image data, where discerning vascular structures impacted by ophthalmological and cardiovascular diseases can be complex.
To diagnose and treat these diseases, the morphological attributes of the retinal vessels are fundamental, including length, width, tortuosity, branching patterns, and angles yao2016convolutional . Although the doctors have the image obtained from medical equipment liu2019retinal , it cannot show the exact structural elements due to imaging equipment limitations and the inherent characteristics of biological tissues. Eventually, the medical experts are required to make the segmentation manually, which is time-consuming and highly prone to bias.
For instance, the most common disease, diabetic retinopathy, which is caused by a higher complication in diabetics and leads to blindness, involves the thickening of the retinal blood vessels’ wall, microvascular occlusions, retinal tissue ischemia, and finally the growth of abnormal new blood vessels, leading to severe complications and the loss of functional vision wang2009retinal .
@@
With the advent of high-performance computing devices and convolutional neural network technology, biomedical image segmentation has shown attractive improvements. In the case of specific architectures with symmetric contracting and expanding paths based on a fully convolutional network, called U-Net ronneberger2015u , a high number of successful research works have been proposed with extension and modification in terms of their general structure. This fully convolutional network architecture has been used in different domains for segmentation tasks kazerouni2021ghost ; xiao2020segmentation ; ICARCV-NIR . In addition, the current transformer networks are not able to reach their performance limit liu2021swin . However, the plain U-Net architecture also has its own limitations, including segmentation of different scale inputs. For a better understanding, we would like to coin a classification of semantic segmentation into two types: small- and large-scale semantic segmentation.
@@
To address these issues, several studies have been conducted, resulting in a number of architectures that have demonstrated improvements, including the salient U-Net hu2019s , structured dropout U-Net guo2019sd , deep residual U-Net liu2023resdo , to name a few. Notwithstanding significant advancements in the extraction of vascular structures, many studies still suffer from microvascular segmentation. Hence, a methodology that can bridge this gap is much desired.
@@ 2 Related Work
Retinal vessel segmentation aims to classify each pixel in a retinal image as either a vessel or background. U-Net ronneberger2015u is commonly used for biomedical image segmentation. To prevent overfitting, methods like regularization and dropout layers are employed jindal2016learning . Spatial information in convolutional layers can lead to overfitting, as noted in SD-U-Net guo2019sd . A structured drop block method is suggested to mitigate this. Training neural networks with dynamic weights was used to overcome problems such as under-segmentation of faint vessels and edge pixels khanal2020dynamic .
Both supervised and unsupervised algorithms are combined for improved segmentation, incorporating multi-scale filters and U-Net models ma2021multichannel . Edge detection and artificial neural networks help in identifying vessel pixels. Hard attention mechanisms with multiple decoder networks target different segmentation regions wang2020hard . Single-step networks can misclassify pixels, leading to the development of NFN+ (Network after Network) with dual multi-scale backbones wu2020nfn+ . Multi-scale dense networks address vessel morphology complexity. Augmentation modules, like channel-wise random Gamma correction, are proposed to handle data variation sun2021robust .
Study group learning (SGL) zhou2021study proposed a method for addressing overfitting in small datasets by automatically erasing vessel segments in the vessel segmentation label. The current state of the art is also U-Net-based, called FR-UNet (Full resolution U-Net). The authors in liu2022full claim that the previous state-of-the-art kamran2021rv is characterized by lower sensitivity metrics, implying that the architecture is not capable of extracting thin vessels efficiently. FR-UNet includes a feature aggregation module that is embedded before each convolution block to aggregate feature maps from up-sampling and down-sampling. This helps extract multi-scale, high-level contextual information. FR-UNet has achieved the highest sensitivity of 0.8316, but there is room for improvement in capturing micro-vessel structures.
@@
Kamran et al. kamran2021rv developed the RV-GAN architecture, a GAN-based model for microvascular localization and segmentation. Despite excellent accuracy, the model’s low sensitivity score suggests it struggles with extracting and localizing thin and branching vascular structures.
@@
HT-Net hu2022ht proposed a future fusion block and a future refinement block to solve the problem of extracting different sizes of vessels. To capture long-range dependencies, HT-Net inserted a transformer with self-attention at the bottom of the network. Similarly, to address multi-scale information and dependencies of vascular structures, MTPA-UNet jiang2022mtpa_unet proposed a combination of convolutional and transformative methods. It includes transformer architecture components that are integrated with an encoder-decoder architecture, demonstrating its promising performance.
@@ 3 Methodology
A new residual block is used instead of the original blocks of U-Net architecture. It is responsible for multi-resolution feature representation by avoiding the vanishing of gradient descent when the depth of the network increases he2016deep . As observed in Figure 0(b), it has a dropout layer to avoid overfitting, and this dropout is included only at the encoder side and not used at the decoder side. The reason is that the decoder part is used for feature up-sampling, and having the dropout layer leads to poor performance of the model.
@@
As shown in Figure 0(a), the left side is an encoder architecture responsible for the feature extraction of the retinal vessels. The primary operations of this block are the convolutional layer, batch normalization layer, max-pooling, and leaky-ReLU activation function. As the network goes deeper, the extraction of the global features will increase while the local features will start to disappear in the process. To achieve this, the skipping connections are used to map these features to the corresponding decoder side to enhance feature visibility, though this cannot fully solve the problem at hand. Notwithstanding that the global features are extracted efficiently when the networks get deeper, thin vessel features will be lost due to the resolution loss at the encoder thus making it impossible to retrieve these features in the decoder kamran2021rv . This is caused by the repeated down-sampling operations. To allow the encoder block to extract both the local and the global features while also retaining the micro-vessel structures, the block shown in Figure 0(b) is introduced. This block is integrated between the consecutive encoder blocks and uses two inputs: from the previous encoder and the encoder booster (except for the first booster, which takes input from the previous encoder and the input to the model), which as a result helps to keep better local and global information than using the standalone encoder blocks. To boost the encoder feature representation ability of micro-vascular structures, atrous convolution chen2017deeplab is used instead of the usual one to be able to control the feature resolutions.
@@
In U-shaped segmentation models ronneberger2015u , the bottleneck block, situated between the encoder and decoder, plays a pivotal role. It diverges from conventional encoders by not using a max-pooling operation and is essential for enhancing feature representation and localization for the decoder. This block compresses the input, retaining key features for high-resolution segmentation map reconstruction.
Our research, drawing inspiration from U-Net3+ huang2020unet , introduces a bottleneck enhancement block. This new component, devoid of the decoder-side dense skip connection to reduce feature redundancy, comprises an up-sampling element and a normalized convolutional layer. It is designed to augment the dense prediction of micro-vessel structures, a task beyond the reach of standard bottlenecks. The block’s integration with the encoder and bottleneck shows promise in enhancing the up-sampling process.
@@
In order to access more contextual information and capture channel-wise dependencies, squeeze and excitation operations are used, respectively hu2018squeeze . Let $F$ be the feature map produced by a convolutional layer. The squeeze operation $S(F)$ is defined as a global average pooling operation which compresses $F$ into a channel descriptor $S(F)=\frac{1}{N}\sum_{i=1}^{N}F_{i}$ where $N$ is the number of pixels in the feature map and $F_{i}$ is the feature value at pixel $i$. The excitation operation is therefore the function $E(S)=\sigma(W_{2}\delta(W_{1}S))$ where $\sigma$ is the sigmoid function, $\delta$ is a ReLU function, and $W_{1},W_{2}$ are the weights of the linear layers.
@@
[…]
@@ 4 Data
Data
Size
Resolution
FoV
Camera
DRIVE staal2004ridge
40
565$\times$584
${}^{\circ}$
Canon CR5 3CCD
CHASE-DB1 fraz2012ensemble
28
990$\times$960
${}^{\circ}$
Nidek NM-200-D
STARE hoover2000locating
20
700$\times$605
${}^{\circ}$
TopCon TRV-50
The DRIVE dataset is divided into 20 training images and 20 testing images. The CHASE-DB1 has 20 training images and 8 testing images. Because the STARE dataset only contains 20 images, k-fold cross-validation is utilized, with k set to 20. All three datasets are independently annotated by two experts, and the one annotated by the first expert is used in this work.
@@ 5.1 Experiment Details
Fig. 2: Comparison of prediction maps of different models with this work on the three datasets.
@@
Table 2: Performance Comparison on DRIVE, CHASE-DB1, and STARE dataset
Dataset
Methods
Sen.
Spe.
F1
Acc.
AUC
IoU
DRIVE
U-Netronneberger2015u
0.7881
0.987
0.8176
0.9695
0.9874
0.6918
Attention UNet oktay2018attention
0.7999
0.986
0.82
0.9695
0.9875
0.6953
U-Net++zhou2018unet++
0.8055
0.9854
0.8208
0.9695
0.9874
0.6964
U-Net3+huang2020unet
0.7874
0.9869
0.8166
0.9693
0.987
0.6905
RV-GANkamran2021rv
0.7927
0.9969 0.8690 0.9790
0.9887
SGLzhou2021study
0.8380
0.9834
0.8316
0.9705
0.9886
SA-UNet guo2021sa
0.8212
0.9840
0.8263
0.9698
0.9864
FR-UNet liu2022full
0.8356
0.9837
0.8316
0.9705
0.9889 0.7120
FS-Net 0.8421
0.9828
0.8314
0.9702
0.9884
0.7116
CHASE-DB1
U-Netronneberger2015u
0.7968
0.9876
0.8038
0.9755
0.9891
0.6723
Attention UNet oktay2018attention
0.7505
0.9908
0.795
0.9756
0.9902
0.6601
U-Net++zhou2018unet++
0.7468
0.9914
0.7961
0.9759
0.9901
0.6618
RV-GANkamran2021rv
0.8199
0.9806
0.8957
0.9697
0.9914
SGLzhou2021study
0.8690
0.9843
0.8271
0.9771 0.9920
FR-UNet liu2022full
0.8798
0.9814
0.8151
0.9748
0.9913
0.6882
FS-Net
0.8365
0.985
0.8144
0.9755
0.9903
0.683
STARE
U-Netronneberger2015u
0.8331
0.9842
0.82
0.9729
0.9903
0.6972
Attention UNet oktay2018attention
0.7813
0.9865
0.7948
0.9714
0.9875
0.6673
U-Net++zhou2018unet++
0.8392
0.9839
0.8231
0.9733
0.9904
0.7009
R2U-Net alom2019recurrent
0.8298
0.9862
0.8475
0.9712
0.9914
U-Net3+huang2020unet
0.8377
0.9813
0.8086
0.9708
0.9791
0.6807
RV-GANkamran2021rv
0.8356
0.9864
0.8323
0.9754
0.9887
FS-Net 0.8393
0.9852
0.8261
0.975
0.9916 0.7055

### v4

@@ Abstract
Retinal vascular segmentation, a widely researched topic in biomedical image processing, aims to reduce the workload of ophthalmologists in treating and detecting retinal disorders. Segmenting retinal vessels presents unique challenges; previous techniques often failed to effectively segment branches and microvascular structures. Recent neural network approaches struggle to balance local and global properties and frequently miss tiny end vessels, hindering the achievement of desired results. To address these issues in retinal vessel segmentation, we propose a comprehensive micro-vessel extraction mechanism based on an encoder-decoder neural network architecture. This network includes residual, encoder booster, bottleneck enhancement, squeeze, and excitation building blocks. These components synergistically enhance feature extraction and improve the prediction accuracy of the segmentation map. Our solution has been evaluated using the DRIVE, CHASE-DB1, and STARE datasets, yielding competitive results compared to previous studies. The AUC and accuracy on the DRIVE dataset are 0.9884 and 0.9702, respectively. For the CHASE-DB1 dataset, these scores are 0.9903 and 0.9755, respectively, and for the STARE dataset, they are 0.9916 and 0.9750. Given its accurate and robust performance, the proposed approach is a solid candidate for being implemented in real-life diagnostic centers and aiding ophthalmologists.
@@
Highlights should contain 3 to 5 bullet points with a maximum of 125 characters per bullet point, including spaces. Only the core results of the paper should be covered. Please ensure each bullet point is less than 125 characters
1. Encoder booster and bottleneck enhancement blocks are introduced to reduce feature loss in the encoder-decoder architecture 2. An adaptive threshold algorithm is proposed to find the optimal threshold for pixelwise classification 3. Micro-retinal vessel extraction was improved significantly 4. A cross-training evaluation was performed between the DRIVE and STARE datasets
@@ 1 Introduction
Medical images are crucial for diagnoses [1], providing valuable information for medical experts to perform necessary procedures. However, these images can be noisy and difficult to interpret, particularly at the microlevel. This challenge is especially noticeable in retinal image data, where discerning vascular structures impacted by ophthalmological and cardiovascular diseases can be complex.
To diagnose and treat these diseases, the morphological attributes of the retinal vessels are fundamental, including length, width, tortuosity, branching patterns, and angles [2]. Although the doctors have the image obtained from medical equipment [3], it cannot show the exact structural elements due to imaging equipment limitations and the inherent characteristics of biological tissues. Eventually, the medical experts are required to make the segmentation manually, which is time-consuming and highly prone to bias.
For instance, the most common disease, diabetic retinopathy, which is caused by a higher complication in diabetics and leads to blindness, involves the thickening of the retinal blood vessels’ wall, microvascular occlusions, retinal tissue ischemia, and finally the growth of abnormal new blood vessels, leading to severe complications and the loss of functional vision [4].
@@
With the advent of high-performance computing devices and convolutional neural network technology, biomedical image segmentation has made significant improvements. In the case of specific architectures with symmetric contracting and expanding paths based on a fully convolutional network, called U-Net [5], a high number of successful research works have been proposed with extension and modification in terms of their general structure. This fully convolutional network architecture has been used in different domains for segmentation tasks [6, 7, 8]. In addition, the current transformer networks are not able to reach their performance limit [9]. However, the plain U-Net architecture also has its own limitations, including segmentation of different scale inputs. For a better understanding, we would like to coin a classification of semantic segmentation into two types: small- and large-scale semantic segmentation.
@@
To address these issues, several studies have been conducted, resulting in a number of architectures that have demonstrated improvements, including the salient U-Net [10], structured dropout U-Net [11], deep residual U-Net [12], to name a few. Notwithstanding significant advancements in the extraction of vascular structures, many studies still suffer from microvascular segmentation. Hence, a methodology that can bridge this gap is much desired.
@@ 2 Related Work
Unsupervised Methods
This branch of machine learning is also applicable in scenarios such as image segmentation, as it has the advantage of not requiring labeled data. Unsupervised methods for image segmentation typically involve three key steps: pre-processing, segmentation, and post-processing. Upadhyay et al. [13] applied two multi-scale transformations—wavelets and curvelets. The former is used to enhance local structures, while the latter is employed to suppress non-vessel backgrounds. Tavakoli et al. [14] introduced the use of random transformations and overlapping windows to detect and locate a vascular tree within a fundus image. Alhussein et al. [15] proposed the application of the Hessian matrix approach to extract both thick and thin vessels, as well as intensity transformation to maximize vessel details. Unsupervised learning relies on pixel colors and intensities while often overlooking contextual information and semantic understanding. In contrast, supervised learning methods such as deep learning yield better outcomes for image data, as they can learn complex features through hierarchical feature learning. This capability enables them to capture detailed information that might be missed by unsupervised methods, such as thin vessels.
@@
Retinal vessel segmentation aims to classify each pixel in a retinal image as either a vessel or background. U-Net [5] is commonly used for biomedical image segmentation. To prevent overfitting, methods like regularization and dropout layers are employed [16]. Spatial information in convolutional layers can lead to overfitting, as noted in SD-U-Net [11]. A structured drop block method is suggested to mitigate this. Training neural networks with dynamic weights was used to overcome problems such as under-segmentation of faint vessels and edge pixels [17].
Both supervised and unsupervised algorithms are combined for improved segmentation, incorporating multi-scale filters and U-Net models [18]. Edge detection and artificial neural networks help in identifying vessel pixels. Hard attention mechanisms with multiple decoder networks target different segmentation regions [19]. Single-step networks can misclassify pixels, leading to the development of NFN+ (Network after Network) with dual multi-scale backbones [20]. Multi-scale dense networks address vessel morphology complexity. Augmentation modules, like channel-wise random Gamma correction, are proposed to handle data variation [21].
Study group learning (SGL) [22] proposed a method for addressing overfitting in small datasets by automatically erasing vessel segments in the vessel segmentation label. The current state of the art is also U-Net-based, called FR-UNet (Full resolution U-Net). The authors in [23] claim that the previous state-of-the-art [24] is characterized by lower sensitivity metrics, implying that the architecture is not capable of extracting thin vessels efficiently. FR-UNet includes a feature aggregation module that is embedded before each convolution block to aggregate feature maps from up-sampling and down-sampling. This helps extract multi-scale, high-level contextual information. FR-UNet has achieved the highest sensitivity of 0.8316, but there is room for improvement in capturing micro-vessel structures.
@@
Kamran et al. [24] developed the RV-GAN architecture, a GAN-based model for microvascular localization and segmentation. Despite excellent accuracy, the model’s low sensitivity score suggests it struggles with extracting and localizing thin and branching vascular structures.
@@
HT-Net [25] proposed a future fusion block and a future refinement block to solve the problem of extracting different sizes of vessels. To capture long-range dependencies, HT-Net inserted a transformer with self-attention at the bottom of the network. Similarly, to address multi-scale information and dependencies of vascular structures, MTPA-UNet [26] proposed a combination of convolutional and transformative methods. It includes transformer architecture components that are integrated with an encoder-decoder architecture, demonstrating its promising performance.
@@ 3 Methodology
A new residual block is used instead of the original blocks of U-Net architecture. It is responsible for multi-resolution feature representation by avoiding the vanishing of gradient descent when the depth of the network increases [27]. As observed in Figure 1(b), it has a dropout layer to avoid overfitting, and this dropout is included only at the encoder side and not used at the decoder side. The reason is that the decoder part is used for feature up-sampling, and having the dropout layer leads to poor performance of the model.
@@
As shown in Figure 1(a), the left side is an encoder architecture responsible for the feature extraction of the retinal vessels. The primary operations of this block are the convolutional layer, batch normalization layer, max-pooling, and leaky-ReLU activation function. As the network goes deeper, the extraction of the global features will increase while the local features will start to disappear in the process. To achieve this, the skipping connections are used to map these features to the corresponding decoder side to enhance feature visibility, though this cannot fully solve the problem at hand. Notwithstanding that the global features are extracted efficiently when the networks get deeper, thin vessel features will be lost due to the resolution loss at the encoder thus making it impossible to retrieve these features in the decoder [24]. This is caused by the repeated down-sampling operations. To allow the encoder block to extract both the local and the global features while also retaining the micro-vessel structures, the block shown in Figure 1(b) is introduced. This block is integrated between the consecutive encoder blocks and uses two inputs: from the previous encoder and the encoder booster (except for the first booster, which takes input from the previous encoder and the input to the model), which as a result helps to keep better local and global information than using the standalone encoder blocks. To boost the encoder feature representation ability of micro-vascular structures, atrous convolution [28] is used instead of the usual one to be able to control the feature resolutions.
@@
In U-shaped segmentation models [5], the bottleneck block, situated between the encoder and decoder, plays a pivotal role. It diverges from conventional encoders by not using a max-pooling operation and is essential for enhancing feature representation and localization for the decoder. This block compresses the input, retaining key features for high-resolution segmentation map reconstruction.
Our research, drawing inspiration from U-Net3+ [29], introduces a bottleneck enhancement block. This new component, devoid of the decoder-side dense skip connection to reduce feature redundancy, comprises an up-sampling element and a normalized convolutional layer. It is designed to augment the dense prediction of micro-vessel structures, a task beyond the reach of standard bottlenecks. The block’s integration with the encoder and bottleneck shows promise in enhancing the up-sampling process.
@@
In order to access more contextual information and capture channel-wise dependencies, squeeze and excitation operations are used, respectively [30]. Let $F$ be the feature map produced by a convolutional layer. The squeeze operation $S(F)$ is defined as a global average pooling operation which compresses $F$ into a channel descriptor $S(F)=\frac{1}{N}\sum_{i=1}^{N}F_{i}$ where $N$ is the number of pixels in the feature map and $F_{i}$ is the feature value at pixel $i$. The excitation operation is therefore the function $E(S)=\sigma(W_{2}\delta(W_{1}S))$ where $\sigma$ is the sigmoid function, $\delta$ is a ReLU function, and $W_{1},W_{2}$ are the weights of the linear layers.
@@
       if $\left|\text{optimum}-\text{ratio}\right|\leq\epsilon$ then
@@ 4 Data
Data Size Resolution FoV Camera
DRIVE [31] 40 565$\times$584 45 Canon CR5 3CCD
CHASE-DB1 [32] 28 990$\times$960 30 Nidek NM-200-D
STARE [33] 20 700$\times$605 35 TopCon TRV-50
DCA1 [34] 134 300$\times$300
The DRIVE dataset is divided into 20 training images and 20 testing images. The CHASE-DB1 has 20 training images and 8 testing images. Because the STARE dataset only contains 20 images, k-fold cross-validation is utilized, with k set to 20. The first three datasets are independently annotated by two experts, and the one annotated by the first expert is used in this work. The DCA1 dataset consists of a total of 134 images with their corresponding ground truth annotated by an expert cardiologist, and it was randomly split into 104 images for training and 30 images for testing.
@@ 5.1 Experiment Details
Fig. 2: Comparison of prediction maps of different models with this work on the Four datasets.
@@
Table 2: Performance Comparison on DRIVE, CHASE-DB1, STARE, and DCA1 datasets.
Dataset Methods Sen. Spe. F1 Acc. AUC IoU
DRIVE Zaho et al.[35] 0.742 0.982 0.954 0.862
U-Net[5] 0.7881 0.987 0.8176 0.9695 0.9874 0.6918
Attention UNet [36] 0.7999 0.986 0.82 0.9695 0.9875 0.6953
U-Net++[37] 0.8055 0.9854 0.8208 0.9695 0.9874 0.6964
U-Net3+[29] 0.7874 0.9869 0.8166 0.9693 0.987 0.6905
Alhussein et al.[15] 0.7851 0.9724 0.9559 0.878
RV-GAN[24] 0.7927 0.9969 0.8690 0.9790 0.9887
SGL[22] 0.8380 0.9834 0.8316 0.9705 0.9886
SA-UNet [38] 0.8212 0.9840 0.8263 0.9698 0.9864
FR-UNet [23] 0.8356 0.9837 0.8316 0.9705 0.9889 0.7120
FS-Net 0.8421 0.9828 0.8314 0.9702 0.9884 0.7116
CHASE-DB1 U-Net[5] 0.7968 0.9876 0.8038 0.9755 0.9891 0.6723
Attention UNet [36] 0.7505 0.9908 0.795 0.9756 0.9902 0.6601
U-Net++[37] 0.7468 0.9914 0.7961 0.9759 0.9901 0.6618
Alhussein et al.[15] 0.7776 0.9634 0.9505 0.8705
RV-GAN[24] 0.8199 0.9806 0.8957 0.9697 0.9914
SGL[22] 0.8690 0.9843 0.8271 0.9771 0.9920
FR-UNet [23] 0.8798 0.9814 0.8151 0.9748 0.9913 0.6882
FS-Net 0.8365 0.985 0.8144 0.9755 0.9903 0.683
STARE U-Net[5] 0.8331 0.9842 0.82 0.9729 0.9903 0.6972
Attention UNet [36] 0.7813 0.9865 0.7948 0.9714 0.9875 0.6673
U-Net++[37] 0.8392 0.9839 0.8231 0.9733 0.9904 0.7009
R2U-Net [39] 0.8298 0.9862 0.8475 0.9712 0.9914
U-Net3+[29] 0.8377 0.9813 0.8086 0.9708 0.9791 0.6807
RV-GAN[24] 0.8356 0.9864 0.8323 0.9754 0.9887
FS-Net 0.8393 0.9852 0.8261 0.975 0.9916 0.7055
DCA1 U-Net[5] 0.7636 0.9895 0.7802 0.9764 0.9908 0.6429
Attention UNet [36] 0.7755 0.9866 0.7687 0.9744 0.9895 0.6276
U-Net++[37] 0.7451 0.9893 0.7661 0.9751 0.9907 0.625
U-Net3+[29] 0.6996 0.9928 0.7611 0.9759 0.9372 0.6196
FS-Net 0.8365 0.9844 0.7909 0.9757 0.9918 0.6559